The Coca-Cola Company and PepsiCo, Inc. provide refreshments to every corner of the world. Suppose selected data from recent consolidated financial statements for The Coca-Cola Company and for PepsiCo, Inc. are presented here (in millions).
Coca-Cola PepsiCo
Total current assets $17,551 $12,571
Total current liabilities 13,721 8,756
Net sales 30,990 43,232
Cost of goods sold 11,088 20,099
Net income 6,824 5,946
Average (net) accounts receivable for the year 3,424 4,654
Average inventories for the year 2,271 2,570
Average total assets 44,595 37,921
Average common stockholders’ equity 22,636 14,556
Average current liabilities 13,355 8,772
Average total liabilities 21,960 23,466
Total assets 48,671 39,848
Total liabilities 23,872 23,044
Income taxes 2,040 2,100
Interest expense 355 397
Net cash provided by operating activities 8,186 6,796
Capital expenditures 1,993 2,128
Cash dividends 3,800 2,732

Collapse question part

(a1)

Correct answer. Your answer is correct.
Compute the following liquidity ratios for Coca-Cola and for PepsiCo. (Round current ratio to 2 decimal places, e.g. 6.25 and all other answers to 1 decimal place, e.g. 15.1.)
Coca-Cola PepsiCo
(1) Current ratio

Entry field with correct answer

: 1

Entry field with correct answer

: 1
(2) Accounts receivable turnover

Entry field with correct answer

times

Entry field with correct answer

times
(3) Average collection period

Entry field with correct answer

days

Entry field with correct answer

days
(4) Inventory turnover

Entry field with correct answer

times

Entry field with correct answer

times
(5) Days in inventory

Entry field with correct answer

days

Entry field with correct answer

days
Click if you would like to Show Work for this question:

Open Show Work

SHOW SOLUTION

LINK TO TEXT

Attempts: 1 of 3 used

Collapse question part

(b1)

Compute the following solvency ratios for the two companies. (Round times interest earned to 1 decimal place, e.g. 15.2, cash debt coverage to 2 decimal places, e.g. 15.25 and percentages to 0 decimal places, e.g. 15%.)
Coca-Cola PepsiCo
(1) Debt to assets ratio

%

%
(2) Times interest earned

times

times
(3) Free cash flow $

$

Click if you would like to Show Work for this question:
Answer #1

a1) Calculation of Liquidity Ratios

Coca-Cola PepsiCo
(1) Current ratio (Total Current Assets/Total Current Liabilities)

Entry field with correct answer(17,551/13,721) = 1.28

: 1

(12,571/8,756) = 1.44Entry field with correct answer

: 1
(2) Accounts receivable turnover (Net Sales/Avg accounts receivable) (30,990/3,424) = 9.1 times

Entry field with correct answer(43,232/4,654) = 9.3

times
(3) Average collection period (365 days/Accounts receivable turnover) (365/9.1) = 40.1 days (365/9.3) = 39.2 days
(4) Inventory turnover (Cost of goods sold/Avg Inventory)

(11,088/Entry field with correct answer2,271) = 4.9

times (20,099/2,570) = 7.8 times
(5) Days in inventory (365 days/Inventory turnover) (365/4.9 times) = 74.5 days (365/7.8 times) = 46.8 days

b1) Calculation of Solvency Ratios

Coca-Cola PepsiCo
(1) Debt to assets ratio (total liabilities/total assets)

[(23,872/48,671)*100] = 49

% [(23,044/39,848)*100] = 58 %
(2) Times interest earned (Earning before interest and tax/Interest expense) [(6,824+2,040+355)/355] = 26.0 times

[(5,946+2,100+397)/397] = 21.3

times
(3) Free cash flow (Net cash provided by operating activities - Capital expenditures) ($8,186-$1,993) = $6,193

($6,796-$2,128) = $4,668

Notes:-

1) Earning before tax and interest = Net Income+Income taxes+interest expense